OpenStudy (anonymous):

Can someone help me how to do this step by step. Given information about one trig function, find other trig functions: if sinθ= 5/8, θ in quadrant II, find sec θ and cot θ.

OpenStudy (anonymous):

find cos theta= 1 - sin^2 (theta) = 1-25/64 from cos , get sec = 1/ cos cot = cos / sin